Gemini Management Systems is a small firm with excellent growth opportunities. Gemini's sales last year were $3 million. They are expected to grow 20% in the coming year. Gemini has no access to any external capital and wants to know whether it can finance its planned growth with its internally generated funds. Gemini requires additional assets equal to 60% of the increase in sales. Short-term liabilities will increase by 20% of the sales increase. The net profit margin is 10%, and Gemini plans to pay a $0.14 million cash dividend next year. Using the AFN equation, will Gemini have enough funds to finance its growth?
Answer #1
A Current Years Sales $3,000,000
B=A*1.2 Expected Sales Next Year $3,600,000
C=B*10% Net Profit $360,000
D Cash dividend $140,000
E=C-D Increase in Retained Earnings $220,000
F=B-A Increase in sales $600,000
G=F*20% Increase in short term liabilities $120,000
H=F*60% Additional assets required $360,000
External Finance Needed=Additional Assets Required -Increase in liabilities -Increase in retained Earnings
I=H-G-E External Finance Needed= $20,000
No , Gemini will not have enough funds to finance growth
It will need external finance = $20,000